# Yogesh Kulkarni, Pooyan Fazli

### v1

@@ Appendix A Appendix
Table 8: Video understanding and reasoning benchmarks: 3B model comparison. AVATAR 3B demonstrates superior performance across both general video understanding and complex reasoning tasks, with particularly strong gains on reasoning-heavy benchmarks like Video-Holmes (+2.9) and MMVU (+2.8). The consistent improvements over baseline GRPO highlight how our stratified replay buffer and TAS enhance learning efficiency even with reduced model capacity, confirming that our framework’s benefits scale effectively across different model sizes. Performance improvements over baseline 3B Omni are shown as subscripts.
@@
Prompt Templates
To support AVATAR’s training pipeline, we use two prompt templates for different components of our framework. The hint generation prompt (Figure 7) is used by our stratified replay buffer’s hint mechanism when the policy becomes stuck in local optima on challenging samples. This prompt instructs a teacher model to generate strategic hints without revealing the final answer, helping the policy escape stagnation while maintaining the learning challenge. The audio-visual localization judge prompt (Figure 8) powers our stepwise reasoning judge in Stage 3 training by providing granular evaluation criteria across four dimensions: grounding of audio cues, identification of visual objects, spatial localization accuracy, and caption correctness. This offers detailed feedback on cross-modal reasoning quality rather than binary success/failure signals.
@@
Figure 8: The prompt template for our VLM-based judge in Stage 3. It guides the judge to produce a granular, stepwise reward by scoring the policy model’s reasoning on audio grounding, visual identification, location accuracy, and caption correctness.
Qualitative Analysis
@@
1def compute_avatar_loss(policy_logprobs: torch.Tensor,
2 old_logprobs: torch.Tensor,
3 advantages: torch.Tensor,
4 completion_mask: torch.Tensor,
5 lambda_tas: float,
6 epsilon: float) -> torch.Tensor:
7
8 # 1. Standard GRPO Clipped Surrogate Objective
9 # ---------------------------------------------
10 # Calculate the importance sampling ratio per token
11 ratio = torch.exp(policy_logprobs - old_logprobs)
12
13 # Clip the ratio to prevent excessively large policy updates
14 clipped_ratio = torch.clamp(ratio, 1 - epsilon, 1 + epsilon)
15
16 # The advantage is broadcasted to match the shape of the per-token ratios
17 advantages_per_token = advantages.unsqueeze(1)
18
19 # Calculate the unclipped and clipped surrogate losses per token
20 loss_unclipped = ratio * advantages_per_token
21 loss_clipped = clipped_ratio * advantages_per_token
22
23 # The standard per-token loss is the minimum of the two, negated
24 per_token_loss = -torch.min(loss_unclipped, loss_clipped)
25
26 # 2. Temporal Advantage Shaping (TAS)
27 # ------------------------------------
28 # Create a tensor to hold the temporal weights, initialized to a baseline of 1.0
29 temporal_weights = torch.ones_like(per_token_loss)
30
31 # Calculate U-shaped weights for each sequence in the batch
32 for i in range(per_token_loss.size(0)):
33 seq_len = completion_mask[i].sum().item()
34 if seq_len > 1:
35 # Normalize token positions to a range of [0, 1]
36 positions = torch.arange(seq_len, device=per_token_loss.device, dtype=torch.float32)
37 normalized_positions = positions / (seq_len - 1)
38
39 # U-Shaped Curve: y = 1.0 + lambda * (2x - 1)^2
40 parabolic_curve = (2 * normalized_positions - 1)**2
41 u_curve_weights = 1.0 + (lambda_tas * parabolic_curve)
42
43 # Apply these weights to the relevant part of the sequence
44 temporal_weights[i, :seq_len] = u_curve_weights
45
46 # 3. Final Loss Calculation
47 # --------------------------
48 # Apply the temporal weights to shape the credit assignment
49 shaped_per_token_loss = per_token_loss * temporal_weights
50
51 # Apply the completion mask to only consider the reasoning tokens
52 masked_loss = shaped_per_token_loss * completion_mask
53
54 # Aggregate the loss for the entire batch
55 final_loss = masked_loss.sum() / completion_mask.sum().clamp(min=1.0)
56
57 return final_loss
Listing 1: Pseudocode for TAS.

### v2

@@ Appendix B Ablation Studies
Table 10: Video understanding and reasoning benchmarks: $3B$ model comparison. AVATAR $3B$ demonstrates superior performance across both general video understanding and complex reasoning tasks, with particularly strong gains on reasoning-heavy benchmarks like Video-Holmes (+2.9+2.9) and MMVU (+2.8+2.8). The consistent improvements over baseline GRPO highlight how our stratified replay buffer and TAS enhance learning efficiency even with reduced model capacity, confirming that our framework’s benefits scale effectively across different model sizes. Performance improvements over baseline 3B Omni are shown in green.
General Video Understanding Video Reasoning Model MVBench Video-MME LVBench Video-Holmes MMVU TOMATO Baseline $3B$ Omni 59.2 57.6 34.3 39.2 62.0 27.0 Baseline GRPO 59.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.6})}$ 58.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.5})}$ 34.7 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.4})}$ 40.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.6})}$ 63.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.2})}$ 27.3 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+0.3})}$ AVATAR $3B$ 61.2 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.0})}$ 59.9 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.3})}$ 35.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.5})}$ 42.1 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.9})}$ 64.8 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+2.8})}$ 28.4 $\scriptstyle{\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}(\textbf{+1.4})}$
B.1 Prompt Templates
To support AVATAR’s training pipeline, we use two prompt templates for different components of our framework. The hint generation prompt (Figure 13) is used by our stratified replay buffer’s hint mechanism when the policy becomes stuck in local optima on challenging samples. This prompt instructs a teacher model to generate strategic hints without revealing the final answer, helping the policy escape stagnation while maintaining the learning challenge. The audio-visual localization judge prompt (Figure 14) helps our stepwise reasoning judge in Stage 33 training by providing granular evaluation criteria across four dimensions: grounding of audio cues, identification of visual objects, spatial localization accuracy, and caption correctness. This offers detailed feedback on cross-modal reasoning quality rather than binary success/failure signals.
GPT-4o Prompt for Critical Step Analysis (TAS Justification) You are an expert evaluator of multimodal reasoning. Your task is to analyze a Chain-of-Thought (CoT) trace and identify the single most critical reasoning step. This is the specific step, phrase, or sentence that, if removed or incorrect, would cause the entire reasoning chain to fail. After identifying this single critical step, you must classify it into one of the following three categories: ‘PLANNING’: The step involves strategy, problem decomposition, or forming an initial hypothesis. (e.g., “I need to first locate the sound, then identify the object.”) ‘EVIDENCE’: The step involves observing unimodal or multimodal facts from the video or audio. (e.g., “The video shows a man wearing a red hat.”, “The audio caption mentions ‘clapping’.”) ‘SYNTHESIS’: The step involves combining multiple pieces of evidence, performing a final calculation, or forming the final conclusion. (e.g., “Since the man is running and the music is fast, the scene is urgent.”, “Count: 3 objects.”) You will be given the user’s query and the model’s reasoning trace. [User Query]: {The original question posed to the model.} [Reasoning CoT]: {The full reasoning trace expected from the model, contained within <think>...</think> tags.} Your Task: Analyze the [Reasoning CoT] and identify the single most critical step. Then, classify it. Constraints: You must output a single, valid JSON object. The JSON object must contain exactly two keys: “category” and “critical_step”. The value for “category” must be one of: PLANNING, EVIDENCE, or SYNTHESIS. The value for “critical_step” must be the exact text of the reasoning step you identified. Example 1 (Planning):
    [User Query]: How many moving rubber objects are there when the video ends?
    [Reasoning CoT]: <think>
    1. I need to identify objects that are ‘rubber’.
    2. I need to check if they are ‘moving’.
    3. I must check this at the ‘end’ of the video.
    4. Then I will count them.
    5. Green cylinder is rubber. It moves from collisions. At the end, it is moving.
    6. Count: 1 moving rubber object.
    </think>
    Output:
    {‘‘category’’: ‘‘PLANNING’’, ‘‘critical_step’’: ‘‘1. I need to identify objects
       that are ‘rubber’. 2. I need to check if they are ‘moving’. 3. I must
       check this at the ‘end’ of the video.’’}
Example 2 (Synthesis):
    [User Query]: What is the audience’s reaction during the woman’s speech?
    [Reasoning CoT]: <think>
    1. The audio caption says ‘‘A crowd cheers and a woman speaks’’.
    2. The video shows a woman at a podium and an audience.
    3. The audio cue ‘‘cheers’’ corresponds to the visual ‘‘audience’’.
    4. ‘‘Cheers’’ are a positive reaction.
    5. Therefore, the audience reaction is positive.
    </think>
    Output:
    {‘‘category’’: ‘‘SYNTHESIS’’, ‘‘critical_step’’: ‘‘5. Therefore, the audience
       reaction is positive.’’}
Figure 12: The GPT-4o [hurst2024gpt] prompt template used to identify and categorize the critical reasoning step from 50,000 SFT samples.
@@ B.1 Prompt Templates
Figure 14: The prompt template for our VLM-based judge in Stage 33. It guides the judge to produce a granular, stepwise reward by scoring the policy model’s reasoning on audio grounding, visual identification, location accuracy, and caption correctness.
B.2 Qualitative Analysis
@@
1def compute_avatar_loss(policy_logprobs: torch.Tensor, 2 old_logprobs: torch.Tensor, 3 advantages: torch.Tensor, 4 completion_mask: torch.Tensor, 5 lambda_tas: float, 6 epsilon: float) -> torch.Tensor: 7 8 # 1. Standard GRPO Clipped Surrogate Objective 9 # --------------------------------------------- 10 # Calculate the importance sampling ratio per token 11 ratio = torch.exp(policy_logprobs - old_logprobs) 12 13 # Clip the ratio to prevent excessively large policy updates 14 clipped_ratio = torch.clamp(ratio, 1 - epsilon, 1 + epsilon) 15 16 # The advantage is broadcasted to match the shape of the per-token ratios 17 advantages_per_token = advantages.unsqueeze(1) 18 19 # Calculate the unclipped and clipped surrogate losses per token 20 loss_unclipped = ratio * advantages_per_token 21 loss_clipped = clipped_ratio * advantages_per_token 22 23 # The standard per-token loss is the minimum of the two, negated 24 per_token_loss = -torch.min(loss_unclipped, loss_clipped) 25 26 # 2. Temporal Advantage Shaping (TAS) 27 # ------------------------------------ 28 # Create a tensor to hold the temporal weights, initialized to a baseline of 1.0 29 temporal_weights = torch.ones_like(per_token_loss) 30 31 # Calculate U-shaped weights for each sequence in the batch 32 for i in range(per_token_loss.size(0)): 33 seq_len = completion_mask[i].sum().item() 34 if seq_len > 1: 35 # Normalize token positions to a range of [0, 1] 36 positions = torch.arange(seq_len, device=per_token_loss.device, dtype=torch.float32) 37 normalized_positions = positions / (seq_len - 1) 38 39 # U-Shaped Curve: y = 1.0 + lambda * (2x - 1)^2 40 parabolic_curve = (2 * normalized_positions - 1)**2 41 u_curve_weights = 1.0 + (lambda_tas * parabolic_curve) 42 43 # Apply these weights to the relevant part of the sequence 44 temporal_weights[i, :seq_len] = u_curve_weights 45 46 # 3. Final Loss Calculation 47 # -------------------------- 48 # Apply the temporal weights to shape the credit assignment 49 shaped_per_token_loss = per_token_loss * temporal_weights 50 51 # Apply the completion mask to only consider the reasoning tokens 52 masked_loss = shaped_per_token_loss * completion_mask 53 54 # Aggregate the loss for the entire batch 55 final_loss = masked_loss.sum() / completion_mask.sum().clamp(min=1.0) 56 57 return final_loss Listing 1: Pseudocode for TAS.